OpenStudy (anonymous):

x+18.4+1/3(54.25-x)=54.25 A very nice person helped me create this equation from a word problem and I feel that I have the right answer but distributing among them is what confuses me. I feel it would be better if I went step by step through this to understand.

OpenStudy (anonymous):

I would feel* sorry about that

jigglypuff314 (jigglypuff314):

(1/3)(54.25-x) = (1/3)(54.25) - (1/3)(x)

OpenStudy (anonymous):

The * is multiplying, yes?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

Also, for 18.4 what happened to it?

OpenStudy (anonymous):

18*4

jigglypuff314 (jigglypuff314):

I was just explaining the distributive property so it was really x + 18.4 + (1/3)(54.25 - x) = 54.25 x + 18.4 + (1/3)*(54.25) - (1/3)*(x) = 54.25

OpenStudy (anonymous):

Oh! That explains it :D

OpenStudy (anonymous):

x+18.4+18.08-1/3x=54.25 The next step is to combine coefficients?

jigglypuff314 (jigglypuff314):

yes :)

jigglypuff314 (jigglypuff314):

x - (1/3)x = (2/3)x so x+18.4+18.08-1/3x=54.25 (2/3)x + 36.48 = 54.25 then subtract 36.48 from both sides - 36.48 -36.48 (2/3)x = 17.77 then multiply both sides by (3/2) *(3/2) *(3/2) x = 17.77*(3/2) then multiply

OpenStudy (anonymous):

26.655?

OpenStudy (anonymous):

I'll check it

jigglypuff314 (jigglypuff314):

yes :) 26.655 was what I got :D

OpenStudy (anonymous):

Thank you again! It felt so good when I got it! You're a lifesaver for this stuff, I wish there was some way I could repay you....

jigglypuff314 (jigglypuff314):

your welcome :) I like helping out with simple algebra :P gives me stuff to procrastinate on lol

OpenStudy (anonymous):

It's amazing how it all fits together when you do it, even though it's not my best suit, but there's always room for improvement >:D Thank you